### v2

# Parallel Key-Value Cache Fusion for Position Invariant RAG

Abstract
Recent advancements in Large Language Models (LLMs) underscore the necessity of Retrieval Augmented Generation (RAG) to leverage external information. However, LLMs are sensitive to the position of relevant information within contexts and tend to generate incorrect responses when such information is placed in the middle, known as the ‘Lost in the Middle’ phenomenon. In this paper, we introduce a framework that generates consistent outputs for decoder-only models, irrespective of the input context order. Experimental results for three open domain question answering tasks demonstrate position invariance, where the model is not sensitive to input context order and superior robustness to irrelevant passages compared to prevailing approaches for RAG pipelines.
Parallel Key-Value Cache Fusion for Position Invariant RAG
Philhoon Oh  Jinwoo Shin  James Thorne KAIST AI {philhoonoh,jinwoos,thorne}@kaist.ac.kr
1 Introduction
In Retrieval Augmented Generation (RAG) (Guu et al., 2020; Lewis et al., 2021; Izacard et al., 2022), models first extract relevant information from a knowledge base and then incorporate this extracted information with its parametric knowledge to generate the response. This two-step approach is the de-facto approach for knowledge-intensive tasks (Lewis et al., 2021; Petroni et al., 2021).
However, decoder-only models exhibit an intrinsic positional bias, assigning more attention to tokens at the beginning or end of the input sequence while often overlooking relevant context located in the middle, a problem known as the ‘Lost in the Middle’ (Liu et al., 2023). Previous works to address this issue involve training with specific prompt (He et al., 2024) or data-intensive training (An et al., 2024). Other works aimed at modifying positional embeddings (Hsieh et al., 2024b) or reducing positional attention bias in LLMs (Yu et al., 2024a). Yet, none of these methods fully guarantee a solution to this intrinsic bias in LLMs for RAG.
Figure 1: Illustration of the KV-Fusion model: Generated tokens remain consistent even when the retrieved passages are shuffled.
Figure 2: Overview of KV-Fusion Architecture. $\mathcal{D}_{\textrm{p}}$ denotes Prefill decoder and $\mathcal{D}_{\textrm{t}}$ represents Trainable decoder. We employ the off-the-shelf LLM to extract the key and value states of the retrieved contexts independently. Then reshaping these caches to train the LLM with task instructions along with questions to generate answers.
In this paper, we introduce a framework for decoder-only models, called Key Value Fusion (KV Fusion), to generate consistent outcomes regardless of input order as illustrated in Figure 1. KV Fusion consists of two components: a prefill decoder that extracts key-value caches in parallel and a trainable decoder that utilizes extracted key-value caches to produce consistent outcomes. This architecture injects uniform positional information into each input context, ensuring consistent output generation even when the input order varies. Experiments on open domain question answering datasets, including NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and POPQA (Mallen et al., 2023), demonstrate KV-Fusion’s position-invariant nature, achieving accuracy improvements of 21.4%, 6.4%, and 6.6% over baseline models in shuffled settings. Furthermore, KV-Fusion models exhibit robust and stable accuracies even with additional contexts compared to other approaches.
2 Method
Notation
Our KV-Fusion architecture is illustrated in Figure 2. For clarity, we refer to this prefill decoder as $\mathcal{D}_{\textrm{p}}$, characterized by the number of key and value heads $|\mathit{H}|$, each with a $\mathit{d}_{h}$ dimension. We denote the trainable decoder as $\mathcal{D}_{\textrm{t}}$, and represent the set of input passages as $\mathcal{C}=\{c_{1},c_{2},\dots,c_{N}\}$ with fixed token length $n$ for each $c_{i}$. This set of passages represents smaller chunks of a long document or retrieved contexts. Lastly, let $L$ represent the total number of layers in $\mathcal{D}_{\textrm{p}}$ and $\mathcal{D}_{\textrm{t}}$, and let $l$ denote the $l$th layer.
Prefill Decoder ($\mathcal{D}_{\textrm{p}}$)
extracts the KV-caches from multiple input passages in parallel, resulting in the injection of identical local positional embeddings ${\{\vec{p_{1}},\vec{p_{2}},\dots,\vec{p_{n}}\}}$. The layer-wise cache representation for each $c_{i}$ is as follows:
$\{k^{l}_{i},v^{l}_{i}\}_{l=1}^{L}=\mathcal{D}_{\textrm{p}}(c_{i}),\quad k^{l}_{i},v^{l}_{i}\in\mathbb{R}^{|\mathit{H}|\times n\times\mathit{d}_{h}}$
Next, we reshape layer-wise KV-caches by concatenating along the token axis over $N$ contexts, forming a single cache for each layer $l$:
$K^{l}=\textsc{res}(\{k^{l}_{i}\}_{i=1}^{N})\quad V^{l}=\textsc{res}(\{v^{l}_{i}\}_{i=1}^{N})$
Here, $K^{l},V^{l}\in\mathbb{R}^{|\mathit{H}|\times(N\times n)\times\mathit{d}_{h}}$ are reshaped KV-cache for the corresponding layer over input passages. These caches prefill and serve as grounding knowledge for training $\mathcal{D}_{\textrm{t}}$.
Trainable Decoder ($\mathcal{D}_{\textrm{t}}$)
takes two inputs: (1) reshaped KV-caches ($\{K^{l},V^{l}\}_{l=1}^{L}$) and (2) target tokens, which contain instruction queries, and answers with a length of $m$ tokens. To ensure sequential alignment of positional information with the KV-caches, position information starting from $\vec{p}_{n+1}$ to $\vec{p}_{n+m}$ are assigned. We then train $\mathcal{D}_{\textrm{t}}$ using next-token prediction, conditioning on the reshaped KV-caches rather than previous tokens:
$\mathcal{D}_{\textrm{t}}(y|q,\mathcal{C^{\prime}})\triangleq\mathcal{D}_{\textrm{t}}(y|q,\{K^{l},V^{l}\}_{l=1}^{L})$
Here, $q$ denotes the instruction with query tokens, and $y$ is answer tokens. $\mathcal{C^{\prime}}$ represents the set of input passages tokens, and $\{K^{l},V^{l}\}_{l=1}^{L}$ is the reshaped KV-caches corresponding to $\mathcal{C^{\prime}}$. We illustrate the details of KV-Fusion in Appendix A.1
3 Experiment Setup
(a) NQ - Llama3
(b) TQA - Llama3
(c) POPQA - Llama3
Figure 3: Comparison of EM Accuracy between KV-Llama3 and Llama3 across different gold context positions. KV-Llama3 maintains its accuracy, while Llama3 shows a tendency for the ‘lost in the middle’ problem.
(a) NQ
(b) TQA
(c) POPQA
Figure 4: Accuracies of baseline and KV models in two scenarios: 1) POS1, where the gold context is positioned first, and 2) Shuffled, where contexts are randomly ordered. KV models maintain their accuracy on both cases, while baseline models struggle in shuffled setting, leading in a wider accuracy gap between the baseline and KV-models.
3.1 Datasets
We consider three open domain question answering datasets: Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and POPQA (Mallen et al., 2023) . [^1] Note that NQ and TriviaQA are filtered version from DPR (Karpukhin et al., 2020) For the base retrieval corpus, we utilize a December 2018 Wikipedia snapshot consisting of 21 million passages, following (Yen et al., 2024; Yu et al., 2024b). Lastly, we use the DPR (Karpukhin et al., 2020) as our baseline retriever to extract the top 40 passages for each dataset.  [^2] Note that we used DPR trained from scratch with its hard negatives on the December 2018 Wikipedia snapshot
Dataset Construction
To enhance the robustness in RAG, we train models with irrelevant contexts (Fang et al., 2024; Yoran et al., 2024a). To this end, we draw the best gold context and extract key phrases among candidate passages by prompting GPT-4o API with a fine-grained template. If all responses are negative, the instance is discarded. Otherwise, we retain the extracted key phrases as evidence, which is later used for training. Negative contexts are sampled from DPR-retrieved passages that do not contain any answer. Each training instance consists of one gold context and 19 negative contexts. The prompt for this process and statistics of all datasets are described in Appendix A.2.
Metric and Evaluations
Exact Match (EM) Accuracy is used for evaluation. (Asai et al., 2023; Mallen et al., 2023). However, we observe that as more documents are added to the input, baseline models tend to generate intrinsic knowledge or hallucinated responses Hsieh et al. (2024a). To address this, we incorporate answerability into the prompt, requiring responses to be concise and limited to a single sentence. Lastly, we set a 48-token limit and use greedy-decoding Huang et al. (2023a) for baselines and KV-Fusions. The template and example for baseline are illustrated in Appendix A.3.
3.2 Training
Input Formatting
Each input passage is formatted with ‘Title:{title}’ and ‘Context:{text}’, followed by a document boundary, ‘====’. For target tokens, we preprend a signal token, <<<|question_answering|>>>, to guide the model’s behavior during inference (Asai et al., 2023). Next, we append instruction and ‘Question:{question}’. Finally, we add answer tokens, which contain both an answer string and a key phrase as evidence, as described in Section 3.1. We hypothesize that appending key phrases enhances the model’s robustness (Thoppilan et al., 2022; Menick et al., 2022). Format examples are provided in Appendix A.4.
Technical Details
We initialize both $\mathcal{D}_{\textrm{p}}$ and $\mathcal{D}_{\textrm{t}}$ with the Llama3-8B model (Dubey et al., 2024). We fine-tune each dataset with a maximum learning rate of $2\times 10^{-5}$ using the AdamW. Across all datasets, we use a batch size of 64 on four A100(80G) GPUs. For the NQ and TQA datasets, models are trained for two epochs. For the POPQA dataset, we fine-tune it on top of the TQA fine-tuned model due to its small training size. The same procedure is applied to the Llama3.1-8B. Detail hyperparameters are reported in Appendix A.5.
4 Results
Dataset NQ TQA POPQA
Top-K 5 10 20 40 5 10 20 40 5 10 20 40
Llama3 34.1 37.2 40.4 38.8 62.4 66.5 67.2 64.7 31.7 33.7 33.7 31.6
Llama3.1 43.2 41.5 42.7 42.4 64.0 64.8 65.9 67.1 31.1 33.4 32.8 33.3
[…]
RePlug-Llama3.1 38.6 36.4 35.3 34.1 65.1 64.4 62.8 60.6 35.4 33.8 30.7 29.4
Pam Qa 51.9 46.5 40.7 19.7 65.9 61.2 52.3 29.0 37.0 35.9 34.9 15.7
KV-Llama3 51.6 51.7 51.4 49.8 67.5 68.8 69.3 69.3 44.5 46.7 48.3 46.7
KV-Llama3.1 51.7 51.8 50.8 49.0 68.6 68.3 69.3 68.7 44.7 47.6 47.4 45.3
Table 1: Percentage accuracy comparison between position-invariant readers. Across top-k results, KV-fusion maintains stable and the strong accuracies, while other models either degrade or exhibit relatively low accuracies.
Model NQ TQA POPQA
Llama3 17.6 14.4 8.2
Llama3.1 14.6 10.6 4.6
KV-Llama3 99.6 99.2 99.4
KV-Llama3.1 99.6 99.8 99.6
Table 2: Percentage Token-Level Match of KV-Fusion and baseline models. While baselines produce varying tokens when contexts are shuffled, KV-Fusion generates identical tokens, even with order perturbations.
Position Invariant RAG
To demonstrate the position-agnostic property, we test models with the gold context placed at varying positions. For each dev dataset, we construct 10 versions by inserting gold context at every alternate location (1st, 3rd, etc.), along with an additional dev set where all 20 contexts are randomly shuffled. To manage the increased inference time, we evaluate the first 500 instances. As shown in Figure 3, KV-Llama3 maintains consistent accuracy across all datasets, regardless of the position of the gold context, while conventional Llama3 shows varying accuracy. A similar pattern is observed with KV-Llama3.1 and Llama3.1 as shown in Appendix A.6. Figure 4 emphasizes this difference: the accuracy of the baseline model drops considerably with shuffled contexts, while the KV models maintain stable performance. In the shuffled scenario, KV-Llama3 achieves higher accuracy than baselines on the NQ, TQA, and POPQA datasets, with similar trends observed for KV-Llama3.1. These findings suggest that KV-Fusion improves performance in RAG.
Comparison with Recent Methods
We evaluate KV models alongside other position-agnostic methods: PAM QA (He et al., 2024), which employs multi-step reasoning to reduce position bias, and RePlug (Shi et al., 2024b), which predicts the next token based on a weighted score for each context. Across the test sets, we utilize up to 40 DPR-retrieved passages, using default settings for PAM QA and the same configurations as Llama3 for RePlug. As shown in Table 1, KV-models achieve the highest accuracy across datasets except the top-5 NQ case. Notably, KV models, originally trained with 20 passages, demonstrate strong robustness even with the top 40 passages. PAM QA performs well with up to 20 passages but shows an average accuracy decline of 50.3% when scaled to the top 40. RePlug follows a similar pattern as the baselines but also experiences performance degradation. Comparable results are observed with contriever passages as shown in Appendix A.7. These results indicate that KV-Fusion enhances robustness even with large input passages within the RAG pipeline.
5 Ablations
Token-Level Consistency
Previous research (He et al., 2024; Hsieh et al., 2024b; Yu et al., 2024a) on positional bias study the effect of performance (e.g. accuracy) under perturbation of the input. Models are regarded as consistent if performance is stable regardless of the position of the gold context. In this section, we further investigate target token-level consistency. We define Token-Level Match (TLM) as the accuracy of Exact Match (EM) between responses in two scenarios: (1) POS1, where the gold context is at the top, and (2) Shuffled, with randomly ordered contexts. Here, $p$ and $s$ denote the responses in POS1 and Shuffled settings, and $N$ is the total number of instances. We experiment with the top-20 dev set in Section 4.
$TLM(p,s,N)=\frac{1}{N}{\sum_{i=1}^{N}EM(p_{i}=s_{i})}$
The results are shown in Table 2. Using greedy-decoding (as this is non-stochastic), Llama3 and Llama3.1 are sensitive to positional context, resulting in low TLM scores across datasets. Notably, the models are most sensitive to positional bias on POPQA. In contrast, KV-Fusion demonstrates nearly perfect TLM scores across all datasets (all $>99\%$). This indicates that KV-Fusion generates near identical tokens regardless of context-order perturbations, highlighting position-agnostic behavior and maintaining token-level consistency where context order is perturbed.
Comparison with Rerankers
To enhance the end-to-end performance of the RAG pipeline, rerankers Nogueira et al. (2020); Zhuang et al. (2023); Ma et al. (2024) are often used to prioritize relevant context while filtering out irrelevant ones (Glass et al., 2022; Yu et al., 2024b). However, we hypothesize that a robust model capable of handling irrelevant passages could eliminate the need for a reranking module. To test this, we rerank the top 40 retrieved passages from the test sets in Section 4 and retain only the top 20 passages, discarding the lower-ranked ones. We utilize point-wise rerankers: MonoT5-3BNogueira et al. (2020), Rankt5-3BZhuang et al. (2023), and RankLLama-7BMa et al. (2024). We then compared the EM accuracy of baseline models evaluated on these top-20 reranked test sets with KV-Fusion models evaluated on the original top-40 test sets.
Results are described in Table 3. Ranking and truncation improve the accuracy of Llama3 across datasets. However, LLama3.1 exhibits slight gains in POPQA but a minor decline in NQ, TQA. This behavior is likely due to LLama3.1’s ability to process longer sequences, making reranking and truncation less beneficial. In contrast, KV-Fusion demonstrates consistently superior accuracy across all datasets, showing its robustness and suggesting that reranking may not be necessary for the RAG.
6 Related Works
Model NQ TQA POPQA
Llama3$\ast$ 38.8 64.7 31.6
[…]
+ rankt5 41.4 67.2 35.4
+ rankllama 40.6 67.4 35.2
Llama3.1$\ast$ 42.4 67.1 33.3
[…]
+ rankt5 42.8 65.9 35.6
+ rankllama 42.0 66.1 35.4
KV-Llama3 $\ast$ 49.8 69.3 46.7
KV-Llama3.1 $\ast$ 49.0 68.7 45.3
Table 3: Accuracy comparison of rerankers based on top-20 reranked contexts, with non-reranked ones using original top-40 contexts ($\ast$). Results show KV-Fusion eliminates the need for ranking in the RAG pipeline.
Retrieval Augmented Generation (RAG)
With recent advancements in LLMs(Team et al., 2024; OpenAI, 2024), Retrieval Augmented Generation (RAG) have proven to be effective in complementing LLMs across various tasks: managing long-tail information (Mallen et al., 2023), reducing hallucinations (Huang et al., 2023b; Shi et al., 2024a), and improving interpretability (Borgeaud et al., 2022; Rudin et al., 2021). The idea of utilizing external knowledge has become prevalent, particularly in knowledge-intensive (Thorne et al., 2018; Lewis et al., 2021; Petroni et al., 2021), where retrievers like DPR and Contriever (Karpukhin et al., 2020; Izacard et al., 2021) first retrieve relevant information, and readers like FiD, ATLAS (Izacard and Grave, 2020; Izacard et al., 2022) incorporate the retrieved information to make predictions.
Robustness and Bias in RAG Pipeline
Despite the promising capabilities of the RAG system, one major challenge is the notable drop in performance when irrelevant contexts exist during inference. (Shi et al., 2023; Oh and Thorne, 2023), along with incorrect responses even when the gold context appears in the middle (Liu et al., 2023). To address these issues, Xu et al. 2023 trained an auxiliary LLM to summarize and extract relevant contexts, while Yoran et al. 2024b proposed a simple Natural Language Inference (NLI) model to eliminate unnecessary passages. Also, He et al. 2024 suggests decomposing inference into multi-step reasoning, enabling the model to generate accurate responses regardless of the context order. Other methods focus on internal features, such as adjusting position hidden states or calibrating attention biases (Hsieh et al., 2024b; Yu et al., 2024a). However, none of these approaches fully resolve a complete solution for the ‘Lost in the Middle’ problem.
Key Value Cache in RAG
Recent studies in Key-Value Cache (KV cache) in RAG focus on improving long context understanding and runtime efficiency. For example, RAGCache Jin et al. (2024) leverages precomputed KV caches for retrieval and faster inference. Similarly, Cache Augmented Generation (CAG) Chan et al. (2024) augments the knowledge via KV caches, demonstrating its efficiency for long-context understanding. TurboRAG Lu et al. (2024) utilizes KV caches for training the RAG pipeline, whereas our work focuses on using KV caches to address the ‘Lost in the Middle’ problem, enhancing robustness and enabling an order-invariant pipeline.
7 Conclusion
This paper presents KV-Fusion, a lightweight training scheme that addresses positional bias and improves the robustness of decoder-only models in the RAG pipeline. KV-Fusion trains language models to be context-order invariant by extracting and pre-filling KV caches with identical positional information and then training decoder-only models using these caches. The results highlight not only the robustness of KV-Fusion in handling a large number of input passages but also its position-invariant property. Our empirical evaluations on three open-domain datasets indicate that KV-Fusion can improve the performance and reliability of the RAG system.
8 Limitations
One limitation of this work is its focus on question answering. Although the most common dataset for evaluating LLMs’ understanding with large context would be the needle-in-a-haystack (NIAH) dataset Kamradt (2023), our experiments are centered around question-answering, which is more challenging than NIAH Hsieh et al. (2024a).
Second limitation is that our experiments are limited to single-hop question answering, where multi-step reasoning is not required. For example, datasets like HotpotQA Yang et al. (2018), and MuSiQue Trivedi et al. (2022) require multiple passages to derive answers. This work, however, focuses on single-hop question-answering datasets, making it difficult to assess the impact of KV-fusion in multi-hop datasets.
Third limitation is that this work does not fully explore the use of KV-cache for training LLMs. Recently, training LLMs by conditioning key-value caches have gained attention Sun et al. (2024) for its efficiency, though our approach remains underexplored in terms of language modeling. However, we present strong empirical results to solve the ‘Lost in the middle’ problem. We hope our work can facilitate future studies on utilizing key-value cache for training LLMs.
Appendix A Appendix
A.1 Details for KV-Fusion Implementation
This section describes the pseudocode for KV-Fusion and Python implementation of the res function.
A.1.1 Algorithm for KV-Fusion
This section further elaborates on the KV-Fusion algorithm, which can be implemented using standard language modeling. For clarification, we provide the pseudocode with a single-instance example. As explained in Section 2, KV-Fusion is built upon two decoders. First, $\mathcal{D}_{\textrm{p}}$ processes a set of input passages retrieved by retrievers, ${C}=\{c_{1},c_{2},\dots,c_{N}\}$, and generates Key-Value (KV) caches in parallel. These KV caches are reshaped by the res function into the form $\{K^{l},V^{l}\}_{l=1}^{L}$ to prefill the cache in $\mathcal{D}_{\textrm{t}}$. Next, $\mathcal{D}_{\textrm{t}}$ processes target tokens, $t=\{t_{1},t_{2},\dots,t_{m}\}$, along with their positional information, $p=\{p_{n+1},p_{n+2},\dots,p_{n+m}\}$. Specifically, the target tokens consist of two parts: the query part, which includes instructions, $q=\{t_{1},t_{2},\dots,t_{k}\}$, and the answer part, $y=\{t_{k+1},t_{k+2},\dots,t_{m}\}$. Along with the prefilled KV cache, we train $\mathcal{D}_{\textrm{t}}$ by prompting it with $q$ and using the standard language model loss to generate $y$. For implementation, we use huggingface transformersWolf et al. (2020) and PyTorchPaszke et al. (2017) libraries.
Algorithm 1 Key Value Fusion(KV Fusion)
Input: $\mathcal{D}_{\textrm{t}}$, $\mathcal{D}_{\textrm{p}}$, Training Data $\mathcal{C^{\mathtt{train}}}=\{C_{1},C_{2},\dots,C_{L}\}$ where ${C_{i}}=\{c^{i}_{1},c^{i}_{2},\dots,c^{i}_{N}\}$, Corresponding query tokens $q_{i}=\{t^{i}_{1},t^{i}_{2},\dots,t^{i}_{k}\}$, Corresponding answer tokens $y_{i}=\{t^{i}_{k+1},t^{i}_{k+2},\dots,t^{i}_{m}\}$
1:Initialize $\mathcal{D}_{\textrm{t}}$, $\mathcal{D}_{\textrm{p}}$ and freeze $\mathcal{D}_{\textrm{p}}$
2:for $i=1,2,...,L$ do
3:     # Extract KV-caches in parallel
4:     ${KV}_{\mathtt{cache}}=\mathcal{D}_{\textrm{p}}(C_{i})$
5:
6:     # Reshape KV-caches
7:     $\{K^{l},V^{l}\}_{l=1}^{L}=\textsc{res}({KV}_{\mathtt{cache}})$
8:
9:     # Compute loss and Optimize $\mathcal{D}_{\textrm{t}}$
10:     $Loss=\mathcal{LM}_{\mathtt{loss}}(y_{i},\mathcal{D}_{\textrm{t}}(q_{i};\{K^{l},V^{l}\}_{l=1}^{L}))$
11:     Update parameters of $\mathcal{D}_{\textrm{t}}$ with respect to $Loss$ via gradient descent
12:end for
A.1.2 res Implementation
To train $\mathcal{D}_{\textrm{t}}$ seamlessly with huggingface transformersWolf et al. (2020) and PyTorchPaszke et al. (2017), extracted KV-cache need to be reshaped to prefill the caches in $\mathcal{D}_{\textrm{t}}$. To this end, we implement res function down below, which can also process batch of instances.
def reshape_key_value_batches(cur_past_key_values, n_psgs):     """     Reshape key-value pairs in batches     """     new_key_cache = []     # Iterate through each key-value pair     for k, v in cur_past_key_values:         # Split keys and values into splits (split by instance)         k_splits, v_splits = torch.split(k, n_psgs, dim=0), torch.split(v, n_psgs, dim=0)         # Reshape and concatenate splits (reshape by instance)         k_re = torch.cat([torch.cat(torch.split(k_val, 1, dim=0), dim=2) for k_val in k_splits], dim=0)         v_re = torch.cat([torch.cat(torch.split(v_val, 1, dim=0), dim=2) for v_val in v_splits], dim=0)         # Append processed key-value pair         new_key_cache.append((k_re, v_re))     return tuple(new_key_cache)
A.2 Dataset Construction
A.2.1 Prompt Template
Prompt Template for Verifying Gold Passages and Supporting Evidence Your task is to find Evidence from a given Document based on a Question and its corresponding Answer. Specifically, the Document contains the Answer for the given Question. Your job is to extract the Evidence from the document. Here are the Question, Document, and Answer. Question: {QUESTION} Document: {PASSAGE} Answer: {ANSWER} Here is how the evidence should be presented: * Evidence - The Evidence should only consist of sentences or paraphrases taken from the given Document. - The Evidence should retain the same format as in the given Document. - The Evidence should inlcude enough information to derive the given Answer from the given Question. - If the provided Document does not contain sufficient information, generate NONE. * Format - DO NOT WRITE ANY GREETING MESSAGES, just write the evidence only. - In front of the evidence, append the word “Evidence:". - Write [END] after you are done. - Here is the Example Format: Evidence: evidence sentences [END] - Do not include “ in the response. Data Generation:
A.2.2 Dataset Statistics
Dataset Training Dev Test
NQ 47,633 3,036 3,610
TQA 34,648 4,288 1,768
POPQA 6,833 1,190 1,267
Table 4: Dataset Statistics for NQ, TQA, and PQA
NQ and TriviaQA are filtered versions provided by Karpukhin et al. 2020 under the CC BY-NC 4.0. We also use POPQA, available from the HuggingFace datasets [^3] https://huggingface.co/datasets/akariasai/PopQA under MIT License. These datasets are English open domain question answering datasets based on the December 2018 Wikipedia snapshot, preprocessed by Karpukhin et al. 2020.
  • NQ dataset: Following the procedure outlined in Section 3.1, we obtained 47,633, 3,036, and 3,610 instances for the training, dev, and test sets, respectively.
  • TQA dataset: To manage the GPT-4o API budget, the original TQA dev set was split in a 2:1 ratio, resulting in newly defined dev and test sets. This produced 34,648, 4,288, and 1,768 instances for the training, dev, and test sets, respectively.
  • POPQA dataset: The original dataset did not include pre-defined training or dev sets. We split the data in an 8:1:1 ratio. After processing with the GPT-4 API, this resulted in 6,833, 1,190, and 1,267 instances for the training, dev, and test sets, respectively
A.3 Baseline Model Template and Example
Baseline Template Title: {TITLE} Context: {TEXT} ===== Title: {TITLE} Context: {TEXT} ===== . . . ===== Title: {TITLE} Context: {TEXT} ===== Strictly based on listed documents (titles and contexts) above, answer the given question clearly and concisely in a single sentence. If none of the documents provide a valid answer, respond with “Unanswerable”. Question: {QUESTION}? ANSWER:
Baseline Example Title: Nobel Prize in Physics Context: Nobel Prize in Physics The Nobel Prize in Physics () is a yearly award given by the Royal Swedish Academy of Sciences for… ===== Title: Nobel Prize Context: His son, George Paget Thomson, received the same prize in 1937 for showing that they also have the properties of waves… ===== . . . ===== Title: Nobel Prize Context: Wilhelm Röntgen’s discovery of X-rays and Philipp Lenard’s work on cathode rays. The Academy of Sciences selected Röntgen… ===== Strictly based on listed documents (titles and contexts) above, answer the given question clearly and concisely in a single sentence. If none of the documents provide a valid answer, respond with “Unanswerable”. Question: who got the first nobel prize in physics? ANSWER:
A.4 KV-Model Input Format Template and Example
The following examples outline the input format template along with a concrete example for $\mathcal{D}_{\textrm{p}}$.
Input Template for $\mathcal{D}_{\textrm{p}}$ Title: {TITLE} Context: {TEXT} =====
Input Example for $\mathcal{D}_{\textrm{p}}$ Title: Does He Love You Context: Does He Love You "Does He Love You" is a song written by Sandy Knox and Billy Stritch, and recorded as a duet by American country music artists Reba McEntire and Linda Davis. It was released in August 1993 as the first single from Reba’s album "Greatest Hits Volume Two". It is one of country music’s several songs about a love triangle. "Does He Love You" was written in 1982 by Billy Stritch. He recorded it with a trio in which he performed at the time, because he wanted a song that could be sung by the other two members =====
The following example outlines the input format template along with a concrete example for $\mathcal{D}_{\textrm{t}}$.
Input Template for $\mathcal{D}_{\textrm{t}}$ <<<|question_answering|>>> Using the provided titles and contexts, answer the given question briefly and provide the supporting sentences as evidence. Question: {QUESTION}? Answer: {ANSWER} [RESULT] Evidence: {EVIDENCE} [END]
Input Example for $\mathcal{D}_{\textrm{t}}$ $<$|question_answering|$>$ Using the provided titles and contexts, answer the given question briefly and provide the supporting sentences as evidence. Question: who sings does he love me with reba? Answer: Linda Davis [RESULT] Evidence: "Does He Love You" is a song written by Sandy Knox and Billy Stritch, and recorded as a duet by American country music artists Reba McEntire and Linda Davis. [END]
A.5 Hyperparameters for training
Training Hyperparameters
flash attention 2 True
target token max length 192
number of input contexts 20
input token max length 192
epochs 2
batch size per gpu 2
gradient accumulation 8
learning rate 2e-5
warmup ratio 0.05
scheduler cosine
optimizer adamW
Table 5: Across all datasets, we utilize four A100 80 GB GPUs, with a batch size of 2 per device and a gradient accumulation of 8, consuming approximately 12 hours (48 GPU hours). The number of training passages is 20, consisting of one gold context and 19 negative contexts. Contexts are tokenized with a maximum length of 192 tokens using left padding; if a context exceeds this limit, tokens are truncated from the left. A random sample of 10,000 contexts from the NQ training set showed that 99.0% of contexts fit within this token limit.(Avg: 145 tokens, Std: 14 tokens) The maximum learning rate is set to $2\times 10^{-5}$, using a linear warmup and cosine decay. The warmup ratio is set to 5%. The AdamW optimizer(paged_adamw_32bit) is used with $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
A.6 KV-Llama3.1 and Llama3.1
(a) NQ - Llama3.1
(b) TQA - Llama3.1
(c) POPQA - Llama3.1
Figure 5: Comparison of EM Accuracy between KV-Llama3.1 and Llama3.1 across different gold context positions. With varying gold context positions, KV-Llama3.1 illustrates consistent accuracies across datasets. However, Llama3.1 suffers from the ‘Lost in the middle’ problem, which can be resolved by KV-Fusion models.
A.7 Position Agnostic reader evaluation on Contriever retrieved passages
Dataset NQ TQA POPQA
Top-K 5 10 20 40 5 10 20 40 5 10 20 40
Llama3 29.6 32.0 36.6 36.1 55.1 59.7 61.8 58.7 37.5 39.1 37.7 36.9
Llama3.1 38.8 36.2 37.7 38.4 58.4 57.6 59.4 60.7 38.8 38.8 38.4 38.8
[…]
RePlug-Llama3.1 35.3 34.9 33.4 32.7 61.0 60.9 60.1 60.4 39.9 34.3 30.8 28.8
Pam Qa 49.9 44.1 38.1 18.9 64.4 57.7 52.7 28.5 51.0 48.8 44.0 24.2
KV-Llama3 49.1 50.6 50.3 49.1 65.1 67.1 68.4 68.3 53.9 56.6 54.5 51.9
KV-Llama3.1 48.9 50.9 50.3 48.9 65.7 66.3 68.8 67.8 53.0 54.8 54.1 51.3
Table 6: Accuracy comparison with other position-invariant methods on contriever-retrieved passages. KV-Fusion models achieve the highest accuracies across datasets except for NQ top-5 case. Consistent with the results observed for DPR-retrieved passages in Table 1, KV-Fusion models show strong robustness to the inclusion of additional passages while other methods experience a decline in performance as more passages are added. Notably, the strong performance on POPQA datasets highlights Contriever’s ability to excel on unseen datasets, which KV-Fusion models effectively leverage. This demonstrates that KV-Fusion models can achieve strong performance on different retrievers.